# Propensity Score Methods for Local Test Score Equating: Stratification and Inverse Probability Weighting

### v1

@@ 1 Introduction
Ensuring fairness in educational testing is a critical issue, especially in the context of large-scale assessments where decisions about examinees are often based on their test scores. The concept of fairness in testing typically revolves around comparability: scores from different forms of the same test should be interchangeable. However, in practice, examinees might take different forms of the test, and ensuring that scores from these forms are comparable is the purpose of test score equating (Kolen and Brennan, , 2014). This becomes particularly challenging in the presence of non-equivalent groups, where examinees are not randomly assigned to different forms, leading to systematic differences in the ability distributions of the groups. Such differences must be accounted for in the equating process to ensure that the resulting scores can be meaningfully compared.
@@
In observed-score equating, the goal is to determine comparable scores between two test forms. Depending on the data collection design, different equating methods are used. Two commonly applied designs are the equivalent groups (EG) design and the nonequivalent groups with anchor test (NEAT) design (Kolen and Brennan, , 2014; von Davier et al., , 2004). In the EG design, two random samples of examinees from the same population take different forms of the test. In contrast, the NEAT design equates the test forms by administering a set of common items, known as the anchor test, to both groups of examinees. The anchor test serves as the link between the two test forms, balancing potential differences between the groups. For further details on these designs, see Kolen and Brennan, (2014), von Davier et al., (2004), and González and Wiberg, (2017).
@@
While the NEAT design is often preferred for equating as it can handle heterogeneous test groups, there are situations where no anchor test is available, particularly in large-scale assessments. For example, the Italian INVALSI test and earlier versions of the Swedish Scholastic Aptitude Test (SweSAT) lacked an anchor test despite having non-equivalent groups (Lyrén and Hambleton, , 2011). In such cases, one possible solution is to use a nonequivalent groups with covariates (NEC) design, where covariates are used to adjust for group differences. Covariates-based equating methods have been explored in both linear and non-linear equating frameworks (Bränberg and Wiberg, , 2011; Wiberg and Bränberg, , 2015). These methods make use of examinees’ background information to balance the groups when anchor items are not available. However, no studies have yet explored the use of covariates for local equating, which is the primary focus of this paper.
@@
Common for all test equating methods is that they aim to fulfill certain equating criteria, such as equity and population invariance. However, many of the methods often struggle in real-world scenarios, and part of the reason for this is that they have not taken full consideration of the equating criteria when defining score equivalence (van der Linden and Wiberg, , 2010). Lord’s equity requirement is particularly important here, as it states that the equated scores from two test forms should be indistinguishable for examinees with the same latent ability $\theta$ (Lord, , 1980). Local equating methods (van der Linden, , 2011; van der Linden and Wiberg, , 2010; Wiberg et al., , 2014) attempt to satisfy this requirement by conditioning the equating process on examinee-level information, such as an anchor test score or other proxies for latent ability. Both linear and non-linear local equating methods have been proposed (Wiberg and van der Linden, , 2011), but those have not yet included the possibility of using covariates.
@@
Using covariates in equating is not a new idea (Kolen, , 1990), and several researchers have explored their use in matching or as complementary information (Cook et al., , 1990; Wright and Dorans, , 1993; Longford, , 2015; Dorans et al., , 2008; Hsu et al., , 2002; Liou et al., , 2001). The propensity score (Rosenbaum and Rubin, , 1983), which is a scalar function of the covariates, has also been applied to test equating in a limited capacity. For example, Livingston et al., (1990) were the first to use propensity scores in test equating, while Paek et al., (2006) used propensity score matching to develop a linking relationship between the PSAT and SAT tests. Further, Sungworn, (2009) used propensity scores based on collateral information to improve poststratification equating. Other researchers, including Moses et al., (2010) and Powers, (2010), examined the potential for propensity scores to reduce equating biases by combining anchor test scores or demographic information. Recently, Wallin and Wiberg, (2019) proposed to use propensity scores in the NEC design, however they did not examine the possibility of using it in local equating.
@@
Despite these developments, the use of propensity scores in local equating has not been explored. This paper aims to fill this gap by introducing two new methods: propensity score stratification and inverse probability weighting (IPW) for local equating. Both of these methods have been widely applied in various areas such epidemiology (Hernán and Robins, , 2006; Austin, , 2008), sociology (Thoemmes and Kim, , 2011; Pais, , 2011), and economics (Huber, , 2015; Vikström, , 2017) to estimate causal effects from observational data. Their popularity stems from their intuitive appeal and their ability to reduce multidimensional covariates to a single scalar summary. Here, these methods are designed to address non-equivalent groups when no anchor test is available, using propensity scores as a proxy for latent ability differences between the test groups to be equated. Propensity score stratification divides examinees into strata based on similar propensity scores, aiming to make the groups balanced in terms of the covariates within each stratum. IPW assigns weights to examinees inversely proportional to their probability of group membership, adjusting for population differences across all score levels. Both methods rely on propensity scores, which represent the conditional probability of group membership based on observed covariates. By using propensity scores as proxies for the unobserved latent ability, we extend local equating to cases where traditional anchor-based methods cannot be applied.
@@ 2.1 Notation and Background
Let $X$ and $Y$ denote observed scores on test forms X and Y, respectively, and let $\theta$ represent the unobserved latent ability influencing responses on both tests. Test score equating aims to find a transformation $\varphi(y)$ that makes scores on form Y comparable to scores on form X. Ideally, this transformation should depend on latent ability $\theta$, resulting in the true ability-specific equating transformation $\varphi_{Y\mid\theta}^{*}(y)$. The fundamental requirement for such a transformation is given by Lord’s equity criterion, which states that the conditional distribution of the equated scores on form Y given ability $\theta$ should be identical to the conditional distribution of scores on form X given $\theta$. For any ability level $\theta$, this criterion is represented as:
@@
Traditional equating methods use a single transformation across the entire population, which averages across ability levels. There will thus always be a compromise made at each ability level for the equating transformation, which inevitably introduces bias. The resulting equated scores will be influenced by the shape of the ability distribution, leading to population-dependent transformations. Local equating addresses these limitations by focusing on the conditional distributions of the test scores given $\theta$ rather than the marginal distributions. This leads to a family of equating transformations, each tailored to a specific level of $\theta$, ensuring that examinees at different ability levels are equated given their ability level. By conditioning on the ability, local equating yields examinee-level transformations, allowing for score comparability that respects differences across ability levels. To define the target transformation, let $F_{X}(x\mid\theta)$ and $F_{Y}(y\mid\theta)$ denote the CDFs for scores $X$ and $Y$, conditional on $\theta$. Local equating defines the transformation as follows:
@@
$\varphi(y;p)=F_{X\mid P=p}^{-1}(F_{Y\mid P=p}(y)),$
where $P$ denotes the proxy variable, and $p$ is a specific proxy value. The validity of using such proxies relies on a surrogate variable assumption, which requires that the proxy captures all relevant information about the relationship between test form assignment and potential outcomes. This can be expressed as:
@@
where $T$ denotes the test form assignment. This condition implies that conditioning on the proxy variable $P$ is equivalent to conditioning on the latent ability $\theta$ for the purpose of removing selection bias in the test form assignment. When this assumption holds, the proxy-based local equating transformation consistently estimates the true ability-specific transformation. However, the quality of this approximation depends critically on how well the chosen proxy satisfies this surrogacy condition.
By conditioning on proxies for latent ability, local equating improves on traditional methods by reducing bias and population dependence. As demonstrated in previous research (van der Linden and Wiberg, , 2010; Wiberg and van der Linden, , 2011), traditional methods like chain equating and poststratification equating can exhibit substantial bias due to their reliance on marginal distributions that do not account for examinee differences in ability. In contrast, local equating maintains Lord’s equity criterion by ensuring that the equating transformation is tailored to the specific ability level of each examinee, thereby producing more accurate scores.
@@ 2.2 Anchor-Based Local Equating
where $T_{i}=0$ if subject $i$ took test form X, $T_{i}=1$ if subject $i$ took form Y, $n_{X|A=a}$ and $n_{Y|A=a}$ are the number of examinees with anchor score $a$ in forms $X$ and $Y$, respectively. The conditional standard deviations for each anchor score $a$ are calculated as:
@@ 3 Propensity Score-Based Equating
When no anchor test is available, but covariates are present, equating may be performed using propensity scores. The propensity score $\pi$ is defined as the probability that an examinee was assigned to form $X$ (or $Y$, whichever is set to the active treatment), conditional on a set of covariates $C_{1},C_{2},\dots,C_{k}$:
@@ 3.1 Propensity Score Stratification
Propensity score stratification is a useful method used in observational studies to reduce bias (Rosenbaum and Rubin, , 1984). This method aims to create subgroups of subjects with similar propensity scores, thereby approximating the balance achieved by randomization in experimental studies. The propensity score serves as a balancing score: when subjects are grouped based on similar propensity scores, the distribution of observed baseline covariates is expected to be similar between treated and untreated subjects within each stratum, mimicking a randomized experiment within these subgroups.
Let $\pi(C_{i})$ denote the propensity score for examinee $i$ with covariates $C_{i}$, estimated via logistic regression. The sample is partitioned into $K$ strata based on the quantiles of the estimated propensity scores, with stratum $k$ denoted by $S_{k}$, where $k=1,\ldots,K$. This quantile-based stratification ensures approximately equal numbers of examinees across strata.
@@
$P(T=1\mid C,S_{k})=P(T=1\mid S_{k}),$
indicating that, conditional on stratum membership, the probability of test form assignment becomes independent of the covariates. When this balance condition holds, examinees within each stratum have similar distributions of covariates regardless of their test form assignment. The choice of the number of strata $K$ involves a bias-variance trade-off. Under suitable regularity conditions, the bias of the stratified estimator decreases with the number of strata at a quadratic rate, suggesting that increasing the number of strata can substantially reduce bias. However, this theoretical advantage must be balanced against the need for sufficient sample size within each stratum to ensure stable estimation of the conditional moments.
@@
where $n_{k,X}$ and $n_{k,Y}$ denote the number of examinees taking forms X and Y in stratum $k$, respectively. Under the balance condition, these within-stratum moments consistently estimate the conditional moments of the test score distributions. The stratum-specific equating transformation is then defined as:
@@ 3.1.1 Assumptions and Properties
1. Positivity: Each examinee must have a non-zero probability of taking either test form, ensuring that $10<\pi(C)<$ 1 for all examinees.
@@
3.2 Inverse Probability Weighting (IPW)
IPW is a statistical technique used in causal inference and observational studies to adjust for confounding and selection bias. The method was first introduced by Horvitz and Thompson, (1952) in the context of survey sampling and has since been widely applied in various fields, including epidemiology, economics, and social sciences. The core idea of IPW is to create a pseudo-population where the distribution of confounding variables is balanced between treatment groups (or, in our case, test forms). This is achieved by assigning weights to individual observations, with the weights being inversely proportional to the probability of receiving the treatment (or taking a particular test form) given the observed covariates.
@@ 3.2 Inverse Probability Weighting (IPW)
$\text{ATE}=E[Y(1)-Y(0)]=E\left[\frac{ZY}{\pi(X)}-\frac{(1-Z)Y}{1-\pi(X)}\right].$ (3)
@@ 3.2.1 IPW for Test Equating
As a second method to use propensity scores, we propose a stratified inverse probability weighting approach for local equating. First, we estimate propensity scores $\pi(C_{i})$ using logistic regression on a set of observed covariates $C_{i}$. The sample is then partitioned into $K$ strata based on the quantiles of the estimated propensity scores. Within each stratum $k$, we define stabilized inverse probability weights:
$w_{ki}=\begin{cases}\frac{p_{k,X}}{\pi(C_{i})}&\text{if }T_{i}=1\\ \frac{1-p_{k,X}}{1-\pi(C_{i})}&\text{if }T_{i}=0\end{cases},$
where $p_{k,X}=P(T_{i}=1|S_{i}=k)$ is the proportion of examinees assigned to form X within stratum $k$, and $S_{i}$ denotes the stratum membership for examinee $i$. To mitigate the impact of extreme weights, we apply symmetric trimming at specified quantiles:
@@
$\hat{\mu}_{k,X}^{w}=\frac{\sum_{i\in S_{k}}w_{ki}^{*}(1-T_{i})X_{i}}{\sum_{i\in S_{k}}w_{ki}^{*}(1-T_{i})},\quad\hat{\mu}_{k,Y}^{w}=\frac{\sum_{i\in S_{k}}w_{ki}^{*}T_{i}Y_{i}}{\sum_{i\in S_{k}}w_{ki}^{*}T_{i}},$
@@
$(\hat{\sigma}_{k,X}^{w})^{2}=\frac{\sum_{i\in S_{k}}w_{ki}^{*}(1-T_{i})(X_{i}-\hat{\mu}_{k,X}^{w})^{2}}{\sum_{i\in S_{k}}w_{ki}^{*}(1-T_{i})},$
$(\hat{\sigma}_{k,Y}^{w})^{2}=\frac{\sum_{i\in S_{k}}w_{ki}^{*}T_{i}(Y_{i}-\hat{\mu}_{k,Y}^{w})^{2}}{\sum_{i\in S_{k}}w_{ki}^{*}T_{i}}.$
@@
$\hat{\varphi}_{k,IPW}(y)=\frac{\hat{\sigma}_{k,X}^{w}}{\hat{\sigma}_{k,Y}^{w}}(y-\hat{\mu}_{k,Y}^{w})+\hat{\mu}_{k,X}^{w}.$
This yields a family of stratum-specific equating functions that transform scores on form Y to the scale of form X, accounting for potential differences in the populations taking each form through both stratification and within-stratum weighting.

### v2

@@ 1 Introduction
Ensuring fairness in educational testing is a critical issue, especially in the context of large-scale assessments where decisions about examinees are often based on their test scores. The concept of fairness in testing typically revolves around comparability: scores from different forms of the same test should be interchangeable. However, in practice, examinees might take different forms of the test, and ensuring that scores from these forms are comparable is the purpose of test score equating (Kolen and Brennan, 2014). This becomes particularly challenging in the presence of non-equivalent groups, where examinees are not randomly assigned to different forms, leading to systematic differences in the ability distributions of the groups. Such differences must be accounted for in the equating process to ensure that the resulting scores can be meaningfully compared.
@@
In observed-score equating, the goal is to determine comparable scores between two test forms. Depending on the data collection design, different equating methods are used. Two commonly applied designs are the equivalent groups (EG) design and the nonequivalent groups with anchor test (NEAT) design (Kolen and Brennan, 2014; von Davier et al., 2004). In the EG design, two random samples of examinees from the same population take different forms of the test. In contrast, the NEAT design equates the test forms by administering a set of common items, known as the anchor test, to both groups of examinees. The anchor test serves as the link between the two test forms, balancing potential differences between the groups. For further details on these designs, see Kolen and Brennan (2014), von Davier et al. (2004), and González and Wiberg (2017).
@@
While the NEAT design is often preferred for equating as it can handle heterogeneous test groups, there are situations where no anchor test is available, particularly in large-scale assessments. For example, the Italian INVALSI test and earlier versions of the Swedish Scholastic Aptitude Test (SweSAT) lacked an anchor test despite having non-equivalent groups (Lyrén and Hambleton, 2011). In such cases, one possible solution is to use a nonequivalent groups with covariates (NEC) design, where covariates are used to adjust for group differences. Covariates-based equating methods have been explored in both linear and non-linear equating frameworks (Bränberg and Wiberg, 2011; Wiberg and Bränberg, 2015). These methods make use of examinees’ background information to balance the groups when anchor items are not available. However, no studies have yet explored the use of covariates for local equating, which is the primary focus of this paper.
@@
Common for all test equating methods is that they aim to fulfill certain equating criteria, such as equity and population invariance. However, many of the methods often struggle in real-world scenarios, and part of the reason for this is that they have not taken full consideration of the equating criteria when defining score equivalence (van der Linden and Wiberg, 2010). Lord’s equity requirement is particularly important here, as it states that the equated scores from two test forms should be indistinguishable for examinees with the same latent ability $\theta$ (Lord, 1980). Local equating methods (van der Linden, 2011; van der Linden and Wiberg, 2010; Wiberg et al., 2014) attempt to satisfy this requirement by conditioning the equating process on examinee-level information, such as an anchor test score or other proxies for latent ability. Both linear and non-linear local equating methods have been proposed (Wiberg and van der Linden, 2011), but those have not yet included the possibility of using covariates.
@@
Using covariates in equating is not a new idea (Kolen, 1990), and several researchers have explored their use in matching or as complementary information (Cook et al., 1990; Wright and Dorans, 1993; Longford, 2015; Dorans et al., 2008; Hsu et al., 2002; Liou et al., 2001). The propensity score (Rosenbaum and Rubin, 1983), which is a scalar function of the covariates, has also been applied to test equating in a limited capacity. For example, Livingston et al. (1990) were the first to use propensity scores in test equating, while Paek et al. (2006) used propensity score matching to develop a linking relationship between the PSAT and SAT tests. Further, Sungworn (2009) used propensity scores based on collateral information to improve poststratification equating. Other researchers, including Moses et al. (2010) and Powers (2010), examined the potential for propensity scores to reduce equating biases by combining anchor test scores or demographic information. Recently, Wallin and Wiberg (2019) proposed to use propensity scores in the NEC design, however they did not examine the possibility of using it in local equating.
@@
Despite these developments, the use of propensity scores in local equating has not been explored. This paper aims to fill this gap by introducing two new methods: propensity score stratification and inverse probability weighting (IPW) for local equating. Both of these methods have been widely applied in various areas such epidemiology (Hernán and Robins, 2006; Austin, 2008), sociology (Thoemmes and Kim, 2011; Pais, 2011), and economics (Huber, 2015; Vikström, 2017) to estimate causal effects from observational data. Their popularity stems from their intuitive appeal and their ability to reduce multidimensional covariates to a single scalar summary. Here, these methods are designed to address non-equivalent groups when no anchor test is available, using propensity scores as a proxy for latent ability differences between the test groups to be equated. Propensity score stratification divides examinees into strata based on similar propensity scores, aiming to make the groups balanced in terms of the covariates within each stratum. IPW assigns weights to examinees inversely proportional to their probability of group membership, adjusting for population differences across all score levels. Both methods rely on propensity scores, which represent the conditional probability of group membership based on observed covariates. By using propensity scores as proxies for the unobserved latent ability, we extend local equating to cases where traditional anchor-based methods cannot be applied.
@@ 2.1 Notation and Background
Let $X$ and $Y$ denote observed scores on test forms $\mathcal{X}$ and $\mathcal{Y}$, respectively, and let $\theta$ represent the unobserved latent ability influencing responses on both tests. Test score equating aims to find a transformation $\varphi(y)$ that makes scores on test form $\mathcal{Y}$ comparable to scores on test form $\mathcal{X}$. Ideally, this transformation should depend on latent ability $\theta$, resulting in the true ability-specific equating transformation $\varphi_{Y\mid\theta}^{*}(y)$. The fundamental requirement for such a transformation is given by Lord’s equity criterion, which states that the conditional distribution of the equated scores on test form $\mathcal{Y}$ given ability $\theta$ should be identical to the conditional distribution of scores on test form $\mathcal{X}$ given $\theta$. For any ability level $\theta$, this criterion is represented as:
@@
Traditional equating methods use a single transformation across the entire population, which averages across ability levels. There will thus always be a compromise made at each ability level for the equating transformation, which inevitably introduces bias. The resulting equated scores will be influenced by the shape of the ability distribution, leading to population-dependent transformations. Local equating addresses these limitations by focusing on the conditional distributions of the test scores given $\theta$ rather than the marginal distributions. This leads to a family of equating transformations, each tailored to a specific level of $\theta$, ensuring that examinees at different ability levels are equated given their ability level. By conditioning on the ability, local equating yields examinee-level transformations, allowing for score comparability that respects differences across ability levels. To define the target transformation, let $F_{X\mid\theta}(x)$ and $F_{Y\mid\theta}(y)$ denote the CDFs for scores $X$ and $Y$, conditional on $\theta$. Local equating defines the transformation as follows:
@@
$\varphi(y;z)=F_{X\mid Z=z}^{-1}(F_{Y\mid Z=z}(y)),$
where $Z$ denotes the proxy variable, and $z$ is a specific proxy value. The validity of using such proxies relies on a surrogate variable assumption, which requires that the proxy captures all relevant information about the relationship between test form assignment and potential outcomes. This can be expressed as:
@@
where $T$ denotes the test form assignment. This condition implies that conditioning on the proxy variable $Z$ is equivalent to conditioning on the latent ability $\theta$ for the purpose of removing selection bias in the test form assignment. When this assumption holds, the proxy-based local equating transformation consistently estimates the true ability-specific transformation. However, the quality of this approximation depends critically on how well the chosen proxy satisfies this surrogacy condition.
By conditioning on proxies for latent ability, local equating improves on traditional methods by reducing bias and population dependence. As demonstrated in previous research (van der Linden and Wiberg, 2010; Wiberg and van der Linden, 2011), traditional methods like chain equating and poststratification equating can exhibit substantial bias due to their reliance on marginal distributions that do not account for examinee differences in ability. In contrast, local equating maintains Lord’s equity criterion by ensuring that the equating transformation is tailored to the specific ability level of each examinee, thereby producing more accurate scores.
@@ 2.2 Anchor-Based Local Equating
where $T_{i}=0$ if subject $i$ took test form $\mathcal{X}$, $T_{i}=1$ if subject $i$ took form $\mathcal{Y}$, $n_{X|A=a}$ and $n_{Y|A=a}$ are the number of examinees with anchor score $a$ in test forms $\mathcal{X}$ and $\mathcal{Y}$, respectively. The conditional standard deviations for each anchor score $a$ are calculated as:
@@ 3 Propensity Score-Based Equating
When no anchor test is available, but covariates are present, equating may be performed using propensity scores. The propensity score $\pi$ is defined as the probability that an examinee was assigned to test form $\mathcal{X}$ (or $\mathcal{Y}$, whichever is set to the active treatment), conditional on a set of covariates $C_{1},C_{2},\dots,C_{k}$:
@@ 3.1 Propensity Score Stratification
Propensity score stratification is a useful method used in observational studies to reduce bias (Rosenbaum and Rubin, 1984). This method aims to create subgroups of subjects with similar propensity scores, thereby approximating the balance achieved by randomization in experimental studies. The propensity score serves as a balancing score: when subjects are grouped based on similar propensity scores, the distribution of observed baseline covariates is expected to be similar between treated and untreated subjects within each stratum, mimicking a randomized experiment within these subgroups.
Let $\pi(\mathbf{C}_{i})$ denote the propensity score for examinee $i$ with covariate vector $\mathbf{C}_{i}=(C_{i1},\cdots,C_{ik})^{\top}$, estimated via logistic regression. The sample is partitioned into $K$ strata based on the quantiles of the estimated propensity scores, with stratum $k$ denoted by $S_{k}$, where $k=1,\ldots,K$. This quantile-based stratification ensures approximately equal numbers of examinees across strata.
@@
$P(T=1\mid\mathbf{C},S_{k})=P(T=1\mid S_{k}),$
indicating that, conditional on stratum membership, the probability of test form assignment becomes independent of the covariates. When this balance condition holds, examinees within each stratum have similar distributions of covariates regardless of their test form assignment. The choice of the number of strata $K$ involves a bias-variance trade-off. Under suitable regularity conditions, the bias of the stratified estimator decreases with the number of strata at a quadratic rate, suggesting that increasing the number of strata can substantially reduce bias (Cochran, 1968; Imbens and Rubin, 2015). However, this theoretical advantage must be balanced against the need for sufficient sample size within each stratum to ensure stable estimation of the conditional moments. In practice we chose $K$ for which the absolute standardized mean differences for all covariates fell below a commonly used threshold while within-stratum sample sizes remained adequate, indicating that the accompanying variance inflation was modest relative to the gain in bias reduction.
@@
where $n_{k,X}$ and $n_{k,Y}$ denote the number of examinees taking forms $\mathcal{X}$ and $\mathcal{Y}$ in stratum $k$, respectively. Under the balance condition, these within-stratum moments consistently estimate the conditional moments of the test score distributions. The stratum-specific equating transformation is then defined as:
@@ 3.1.1 Assumptions and Properties
1. Positivity: Each examinee must have a non-zero probability of taking either test form, ensuring that $0<\pi(C)<1$ for all examinees.
@@
3.2 Inverse Probability Weighting
IPW is a statistical technique used in causal inference and observational studies to adjust for confounding and selection bias. The method was first introduced by Horvitz and Thompson (1952) in the context of survey sampling and has since been widely applied in various fields, including epidemiology, economics, and social sciences. The core idea of IPW is to create a pseudo-population where the distribution of confounding variables is balanced between treatment groups (or, in our case, test forms). This is achieved by assigning weights to individual observations, with the weights being inversely proportional to the probability of receiving the treatment (or taking a particular test form) given the observed covariates.
@@ 3.2 Inverse Probability Weighting
$\text{ATE}=E[Y(1)-Y(0)]=E\left[\frac{TY}{\pi(\mathbf{C})}-\frac{(1-T)Y}{1-\pi(\mathbf{C})}\right].$ (3)
@@ 3.2.1 IPW for Test Equating
As a second method to use propensity scores, we propose a stratified inverse‐probability‐weighting approach for local equating. First, we estimate propensity scores $\pi(\mathbf{C}_{i})$ via logistic regression on a set of observed covariates $\mathbf{C}_{i}$. We then partition the sample into $K$ strata according to the quantiles of the estimated $\pi(\mathbf{C}_{i})$. Within each stratum $k$, we define stabilized inverse‐probability weights as
$w_{ki}=\begin{cases}\dfrac{p_{k,Y}}{\pi(\mathbf{C}_{i})},&T_{i}=1,\\[9.24994pt] \dfrac{1-p_{k,Y}}{1-\pi(\mathbf{C}_{i})},&T_{i}=0,\end{cases}$
where
$p_{k,Y}=P\bigl(T_{i}=1\mid S_{i}=k\bigr)$
is the proportion of examinees in stratum $k$ who took form $Y$ (recall $T_{i}=1$ denotes form $Y$ and $T_{i}=0$ denotes form $X$), and $S_{i}$ denotes the stratum membership for examinee $i$. To mitigate the impact of extreme weights, we apply symmetric truncation [^1] Because the propensity-score stratification procedure already limits the influence of extreme $\hat{\pi}(\mathbf{C}_{i})$ by pooling similar scores into homogeneous strata, no additional weight truncation is required there; the two methods are therefore comparable with respect to the treatment of extreme propensities. at specified quantiles:
@@
$\hat{\mu}_{k,\mathcal{X}}^{w}=\frac{\sum_{i\in S_{k}}w_{ki}^{*}(1-T_{i})X_{i}}{\sum_{i\in S_{k}}w_{ki}^{*}(1-T_{i})},\quad\hat{\mu}_{k,\mathcal{Y}}^{w}=\frac{\sum_{i\in S_{k}}w_{ki}^{*}T_{i}Y_{i}}{\sum_{i\in S_{k}}w_{ki}^{*}T_{i}},$
@@
$(\hat{\sigma}_{k,\mathcal{X}}^{w})^{2}=\frac{\sum_{i\in S_{k}}w_{ki}^{*}(1-T_{i})(X_{i}-\hat{\mu}_{k,X}^{w})^{2}}{\sum_{i\in S_{k}}w_{ki}^{*}(1-T_{i})},$
$(\hat{\sigma}_{k,\mathcal{Y}}^{w})^{2}=\frac{\sum_{i\in S_{k}}w_{ki}^{*}T_{i}(Y_{i}-\hat{\mu}_{k,\mathcal{Y}}^{w})^{2}}{\sum_{i\in S_{k}}w_{ki}^{*}T_{i}}.$
@@
$\hat{\varphi}_{k,IPW}(y)=\frac{\hat{\sigma}_{k,\mathcal{X}}^{w}}{\hat{\sigma}_{k,\mathcal{Y}}^{w}}(y-\hat{\mu}_{k,\mathcal{Y}}^{w})+\hat{\mu}_{k,\mathcal{X}}^{w}.$
This yields a family of stratum-specific equating functions that transform scores on test form $\mathcal{Y}$ to the scale of test form $\mathcal{X}$, accounting for potential differences in the populations taking each test form through both stratification and within-stratum weighting.